OpenStudy (kaylala):

topic: PROVING IDENTITIES - i need help please :) i dont know how to do this (SEE COMMENTS) (trigonometry)

OpenStudy (kaylala):

1) [sec s] - [(cos s)/(1 + sin s)] = tan s 2) [(sec A + csc A cot A)/(sec^2 A)] = cos A csc^2 A

OpenStudy (loser66):

2/ ( don't know why, I always start from #2) break down the fraction from the left hand side, what do you have?

OpenStudy (loser66):

\[\frac{sec}{sec^2}+\frac{csc* cot}{sec^2}\] \[=\Large cos +\frac{\frac{1}{sin}*\frac{cos}{sin}}{\frac{1}{cos^2}}\] \[=\Large cos +\dfrac{\dfrac{cos}{sin^2}}{\dfrac{1}{cos^2}}\] \[=\Large cos +cos \frac{cos^2}{sin^2}\\=cos +cos *cot^2\] factor cos out cos (1+cot^2) = cos *csc^2

OpenStudy (kaylala):

okay then. we start with 2 i dont really know. sorry

OpenStudy (loser66):

question?

OpenStudy (kaylala):

how did you get the very first one? @Loser66

OpenStudy (loser66):

break the factor out.